# On difference sets of dense subsets of $\mathbb{Z}^{2}$

### v1

@@ Abstract
Although this question remains open, we establish a relatively weaker form of this conjecture. Specifically, we prove that there exist infinitely many integers $k\in\mathbb{Z}$ and a sequence $\langle x_{n}\rangle_{n\in\mathbb{N}}$ in $\mathbb{Z}$ such that
$k\cdot FS(\langle x_{n}\rangle_{n\in\mathbb{N}})\subseteq\{\,xy:(x,y)\in E-E\,\},$
where $FS(\langle x_{n}\rangle)$ denotes the finite sums set generated by the sequence.
@@ 1  Introduction
The notions of $IP$ sets have significant importance in Ramsey theory, for ex. see [6]. In this article, we prove the following theorem, which is a weaker version of the Problem 1.3.
@@
$k\cdot FS(\langle x_{n}\rangle_{n})\subseteq\{\,xy:(x,y)\in E-E\,\}.$
If for any sequence $\langle x_{n}\rangle_{n}$ in $\mathbb{Z}$ and $H\in\mathcal{P}_{f}\left(\mathbb{N}\right),$ define $x_{H}=\sum_{t\in H}x_{t}.$ Our technique also proves the following result.
Theorem 1.5.
$\{x_{K}-x_{H}:\max H<\min K\}\subseteq\{\,x+y:(x,y)\in E-E\,\}.$
@@ 2  Preliminaries
Note that Central sets can be defined as members of minimal idempotent ultrafilters, and Central* sets are those sets that sit in every minimal idempotent ultrafilter. We will recall these notions in the coming section.
The following theorem, which follows from [1, Theorem 1.4], shows that the set of return times is large.
@@ 3  Our Results
To obtain Theorem 1.4 and Theorem 1.5, we will study the properties of the set $AA^{-1}$, where $A\subseteq G\times G$ with positive upper Banach density. First, we prove a sequence of lemmas. Although we don’t need the following lemma in its full strength, we prove it here for its own interest.
@@
In fact, the above proof suggests something more: if $p\in\mathcal{LDR}_{2}(G)$ and $q\in\mathcal{RDR}_{2}(G),$ then $p\otimes q\in T.$ As $E(K(\beta G,\cdot))\subseteq\mathcal{DR}_{2}(G),$ from the lemma 3.4 the following lemma follows.
@@
Let $E\subseteq\mathbb{Z}\times\mathbb{Z}$ be a subset with positive upper Banach density, then for any $p,q\in E(K(\beta\mathbb{Z},+))$ we have $(E-E)\in p\otimes q.$
@@
Now choose any $p,q\in E(K(\beta\mathbb{Z},+))$. But from our claim, it follows that $f(E-E)\in\beta f(p\otimes q)=p\cdot q.$ Hence $\{\,xy:(x,y)\in E-E\,\}\in p\cdot q.$ As $q\in E(K(\beta\mathbb{Z},+)),$ it is an routine exercise to check that sequences $\langle x_{n}\rangle_{n}$ in $\mathbb{Z}$ such that
$k\cdot FS(\langle x_{n}\rangle_{n})\subseteq\{\,xy:(x,y)\in E-E\,\}.$
@@
$\beta f(E-E)=\{\,x+y:(x,y)\in E-E\,\}\in(-p)+p.$
So to prove our result, it will be sufficient to show that all the members of (p)+p(-p)+p contain the desired pattern. Let $A\in(-p)+p.$ Then $\{x:-x+A\in p\}\in(-p).$ Hence $B=\{y:y+A\in p\}\in p.$ Note $B^{\star}=\{y:-y+B\in p\}\in p.$ Pick $y_{1}\in B^{\star}.$ From [6, Lemma 4.14], $y_{1}+B^{\star}\in p.$ Then pick

### v2

@@ Abstract
Although this question remains open, we establish a relatively weaker form of this conjecture. Specifically, we prove that if $\langle a_{j}\rangle_{j=1}^{m}$ is any finite sequence in $\mathbb{N},$ then there exist infinitely many integers $k\in\mathbb{Z}$ and a sequence $\langle x_{n}\rangle_{n\in\mathbb{N}}$ in $\mathbb{Z}$ such that
$k\cdot MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)\subseteq\{\,xy:(x,y)\in E-E\,\},$
where $MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)$ denotes the milliken-Taylor configuration generated by the sequences $\langle a_{j}\rangle_{j=1}^{m}$ and $\langle x_{n}\rangle_{n\in\mathbb{N}}$.
@@ 1  Introduction
Over the semigroup $(\mathbb{N},\cdot),$ we denote it by $FP\left(\langle x_{n}\rangle_{n\in\mathbb{N}}\right).$ The notions of $IP$ sets have significant importance in Ramsey theory, for ex. it is known that they are partition regular (see [6]).
An extension of Hindman’s Finite Sum Theorem was studied independently by Milliken and Taylor in [8, 10]. To address their results, we need the following notions of Milliken-Taylor system.
Definition 1.4 (Milliken-Taylor system).
Let $m\in\mathbb{N},$ let $\langle a_{j}\rangle_{j=1}^{m}$ and $\langle x_{n}\rangle_{n}$ be sequences in $\mathbb{N}.$ The Milliken-Taylor System determined by $\langle a_{j}\rangle_{j=1}^{m}$ and $\langle x_{n}\rangle_{n}$ is
$MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)=\left\{\sum_{j=1}^{m}a_{j}\cdot\sum_{t\in F_{j}}x_{t}:F_{1}<F_{2}<\cdots<F_{m},\forall F_{i}\in\mathcal{P}_{f}\left(\mathbb{N}\right)\right\}.$
So clearly, for any sequence $\langle x_{n}\rangle_{n},$ $FS(\langle x_{n}\rangle_{n})=MT\left(\{1\},\langle x_{n}\rangle_{n}\right).$ However, the following theorem says a more general result than the Finite Sum Theorem. Although in the original text the result was described over the set of positive integers, the technique for $\mathbb{Z}$ is similar.
Theorem 1.5.
[7, Theorem 17.31] Let $\langle a_{j}\rangle_{j=1}^{m}$ be sequences in $\mathbb{N}.$ Then for any $p\in E(\beta\mathbb{Z},+)$ and any $A\in a_{1}p+\cdots+a_{m}p,$ there exists a sequence $\langle x_{n}\rangle_{n}$ such that $MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)\subseteq A.$
In this article, we prove the following theorem, which is a weaker version of the Problem 1.3.
Theorem 1.6.
Let $m\in\mathbb{N},$ and $\langle a_{j}\rangle_{j=1}^{m}$ be a sequence in $\mathbb{N}$ is given. If $E\subset\mathbb{Z}^{2}$ is a set of positive upper Banach density, then there exist infinitely many $k\in\mathbb{Z}$ and infinitely many sequences $\langle x_{n}\rangle_{n}$ in $\mathbb{Z}$ such that
$k\cdot MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)\subseteq\{\,xy:(x,y)\in E-E\,\}.$
For any sequence $\langle x_{n}\rangle_{n}$ in $\mathbb{Z}$ and $H\in\mathcal{P}_{f}\left(\mathbb{N}\right),$ define $x_{H}=\sum_{t\in H}x_{t}.$ Our technique also proves the following result.
@@
$\{x_{K}-x_{H}:H,K\in\mathcal{P}_{f}\left(\mathbb{N}\right),\text{ and }\max H<\min K\}\subseteq\{\,x+y:(x,y)\in E-E\,\}.$
@@ 2  Preliminaries
Note that Central sets can be defined as members of minimal idempotent ultrafilters, and Central* sets are those sets that sit in every minimal idempotent ultrafilter. We will recall these notions in the coming section The following theorem, which follows from [1, Theorem 1.4], shows that the set of return times is large.
@@ 3  Our Results
To obtain Theorem 1.6 and Theorem 1.7, we will study the properties of the set $AA^{-1}$, where $A\subseteq G\times G$ with positive upper Banach density. First, we prove a sequence of lemmas.
@@
The above proof suggests something more: if $p\in\mathcal{LDR}_{2}(G)$ and $q\in\mathcal{RDR}_{2}(G),$ then $p\otimes q\in T.$
That means if $E\subseteq\mathbb{Z}\times\mathbb{Z}$ is a subset with positive upper Banach density, then for any $p\in\mathcal{LDR}_{2}(G)$ and $q\in\mathcal{RDR}_{2}(G),$ we have $(E-E)\in p\otimes q.$
To prove Theorem 1.6, we need the following lemma. Again, the reference contains the proof over the set of positive integers, but the same proof holds over the set of all integers.
Lemma 3.6.
[7, Lemma 5.19.2] If $p\in E(K(\beta\mathbb{Z},+))$ and $n\in\mathbb{N},$ then $n\cdot p\in E(K(\beta\mathbb{Z},+)).$
@@
Now choose any $p,q\in E(K(\beta\mathbb{Z},+))$. Then from Lemmas 3.6, 3.2 we have $a_{1}q+\cdots+a_{m}q\in\mathcal{RDR}_{2}(\mathbb{Z}).$ But from Lemma 3.5 and our claim, it follows that
$f(E-E)\in\beta f(p\otimes(a_{1}q+\cdots+a_{m}q))=p\cdot(a_{1}q+\cdots+a_{m}q).$
Hence
$\{\,xy:(x,y)\in E-E\,\}\in p\cdot(a_{1}q+\cdots+a_{m}q).$
Now from Theorem 1.5, it is an routine exercise to check that sequences $\langle x_{n}\rangle_{n}$ in $\mathbb{Z}$ such that
$k\cdot MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)\subseteq\{\,xy:(x,y)\in E-E\,\}.$
@@
$f(E-E)=\{\,x+y:(x,y)\in E-E\,\}\in(-p)+p.$
So to prove our result, it will be sufficient to show that all the members of (p)+p(-p)+p contain the desired pattern. Let $A\in(-p)+p.$ Then $\{x:-x+A\in p\}\in(-p).$ Hence $B=\{y:y+A\in p\}\in p.$ Note $B^{\star}=\{y:-y+B\in p\}\in p.$ Pick $y_{1}\in B^{\star}.$ From [7, Lemma 4.14], $y_{1}+B^{\star}\in p.$ Then pick